OpenStudy (anonymous):

Find an equation in standard form of the line containing the given points. (3,-2), (-3,2)

OpenStudy (shamil98):

standard form is ax +by = c first find the slope of the two coordinates and then put it into point-slope form, and then change it to slope-intercept form from there. Then finally change it to standard form.

OpenStudy (shamil98):

2+2 / -3 -3 = 4/-6 = -2/3 Point-slope form is y -y1 = m(x-x1) slope-intercept form is y = mx + b